OpenStudy (anonymous):

how would i solve this.... 1-3y/5 =2

OpenStudy (asnaseer):

is this:\[\text{a) }1-\frac{3y}{5}=2\]or,\[\text{b) }\frac{1-3y}{5}=2\]

OpenStudy (anonymous):

b

OpenStudy (asnaseer):

ok, first step would be to multiply both sides by 5 to get rid of the denominator on the left hand side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer y =- 1/3

OpenStudy (asnaseer):

so you woud get:\[5*\frac{1-3y}{5}=5*2\]\[\cancel{5}*\frac{1-3y}{\cancel{5}}=5*2\]\[1-3y=10\]

OpenStudy (anonymous):

oh okay i found out what i did wrong i multiply 5 with the other numbers on top too.. thank you

OpenStudy (asnaseer):

yw